OpenStudy (anonymous):

Square root of 72x^2?

OpenStudy (anonymous):

i dont knw how to. explain please?

OpenStudy (anonymous):

Do you know what a perfect square is ?

OpenStudy (anonymous):

yes, Perfect square is a number that has a rational number as its square root.

OpenStudy (anonymous):

\[\sqrt{(72x)^{2}}\] Is this the equation you want to be solved?

OpenStudy (anonymous):

If so, \[\sqrt{(72x)^{2}}= 72x\]

OpenStudy (anonymous):

Or is it √72x^2 without the parentheses

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Then you need to find a factor of 72 that is a perfect square.

OpenStudy (anonymous):

In other words, \[\sqrt{72}= 6\sqrt{2}\] So, \[6\sqrt{2}\approx 8.49\] Therefore \[\sqrt{72x ^{2}}=8.49x\]

OpenStudy (anonymous):

Actually, \[\sqrt{72x ^{2}}\approx 8.49x\]